Ticket Number: HEA-NOC/20060815-8 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20060815 13:18 UTC+1 Problem Start: 20060815 12:39 UTC+1
Ticket Update: 20060815 18:51 UTC+1 Problem End:
Site/line: NCAD, NUIM
Link flaps on NUIM (HEAN0005) and NCAD (NCAD103073) BT circuits
Traffic to and from NCAD and NUIM was affected. NUIM's traffic went over
their backup ISDN line during the outage.
20060815 13:09 OMG Called BT support (spoke to Sean) and asked to check
for any unscheduled work being done on these circuits and an alarm check
from 12:20-13:00. Awaiting call back from an engineer. There was PEW on
these circuits yesterday evening (see ticket #20060811-2). BT's ticket
reference is 1-346240587. Both circuits came back at 12:42.
20060815 16:38 OMG Rang BT (Luke) for an update on the alarm check that I
requested. The engineer checked and said that there was an alarm in the
PoP at 12:28 that cleared at 12:31. I asked him to find out the reason for
the alarm, and also to check if there have been further issues with these
circuits as NUIM have reported fluctuations on their link.
20060815 18:50 BS BT called, have not confirmed any other alarms since
12:31, however visibility limited at present due to temporary problem with
This ticket can be monitored at http://www.hea.net/tickets/20060815-8